### v2

# Data-dependent and Oracle Bounds on Forgetting in Continual Learning

Abstract
In continual learning, knowledge must be preserved and re-used between tasks, maintaining good transfer to future tasks and minimizing forgetting of previously learned ones. While several practical algorithms have been devised for this setting, there have been few theoretical works aiming to quantify and bound the degree of Forgetting in general settings. We provide both data-dependent and oracle upper bounds that apply regardless of model and algorithm choice, as well as bounds for Gibbs posteriors. We derive an algorithm based on our bounds and demonstrate empirically that our approach yields tight bounds on forgetting for several continual learning problems.
1 Introduction
Continual learning is a burgeoning machine learning setting where data from different tasks are presented sequentially to the learner. The usual stated goal of methods in this setting is to adapt the learner to new tasks as they appear while also preserving its performance on previous tasks (De Lange et al., 2021; Hadsell et al., 2020; Parisi et al., 2019). This performance on previous tasks is called backward transfer, or forgetting, and one of the key challenges in continual learning is avoiding Catastrophic Forgetting (Goodfellow et al., 2014; Ramasesh et al., 2020; Kirkpatrick et al., 2017), meaning that performance on previous tasks degrades significantly as the model adapts to new tasks.
Although avoiding catastrophic forgetting is desirable, much of the focus of continual learning research in recent years was on settings with a shared optimal solution. Realistically, we should not expect a single algorithm to perform optimally across all settings, for example if the data distribution changes gradually. A recent paper by Kumar et al. (2023) discusses continual learning as a computationally constrained optimization problem and argues that forgetting non-recurring information is not “catastrophic”, especially given changing environments. We will discuss this topic further in our empirical evaluation.
While there have been several empirical approaches in the field, there are relatively few theoretical works that explore and attempt to quantify and bound this backward transfer. Some, such as Evron et al. (2022); Lin et al. (2023); Li et al. (2024) focus on linear models to consider the effect of task order and similarity on forgetting. Others, such as Bennani and Sugiyama (2020); Doan et al. (2021) utilize the NTK regime to focus on more complex task similarity measures as predictors of forgetting. Several more general works such as Benavides-Prado and Riddle (2022) apply notions of VC-dimension to arrive at more general scaling laws and upper bounds on forgetting for the ERM model, but may be difficult to apply for complex predictors due to the potentially large VC-dimension of models such as deep neural networks (Bartlett et al., 2019). We note that many of the known results (Bennani and Sugiyama, 2020; Evron et al., 2022; Benavides-Prado and Riddle, 2022) provide upper bound on forgetting for the training data. Our work, however, focuses on general bounds on forgetting for test data. We note that there are some previous results on forgetting for test data (Doan et al., 2021; Lin et al., 2023; Li et al., 2024) for specific model. In particular, Lin et al. (2023) provide an exact characterization of forgetting under specific assumptions in the linear regression setting.
In this work, we explore upper bounds on forgetting that apply for both general and specific models. We use the PAC-Bayes (McAllester, 1999; Catoni, 2004; Alquier, 2024) framework to derive and analyze upper bounds on backward transfer, focusing on the Gibbs posterior (Casella and George, 1992). In Section 3 we derive general bounds for the two-task and $T$-task settings without making any model assumptions. We then focus our discussion in Section 4 on oracle bounds for the Gibbs posterior in general, and under specific assumptions of task similarity. We then proceed to derive an algorithm based on our bounds and we empirically validate it [^1] Anonymized Code is available in a separate zip file. on several continual learning problems in Section 5. To the best of our knowledge, our bounds are the first general PAC-Bayes upper bounds on forgetting in continual learning that are applicable regardless of model choice and task distribution.
2 Problem definition
We consider a finite sequence of tasks $\{1,2,\ldots,T\}\triangleq[T]$, where for each task $k\in[T]$, we are given a batch of i.i.d. data $S_{k}\sim\mathcal{D}_{k}$. The sample for a given task is defined as $S=\{z_{i}\}_{i=1}^{m},z_{i}=(x_{i},y_{i})$ where $x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y}$. A hypothesis $h\in\mathcal{H}$ is a mapping $h:\mathcal{X}\rightarrow\mathcal{Y}$ characterized by a loss $\ell(h,z)$.
Definition 1.
The expected loss of a given hypothesis $h\in\mathcal{H}$ is defined as $\mathcal{L}(h,\mathcal{D})\triangleq E_{z\in\mathcal{D}}\ell(h,z)$. The empirical loss of a hypothesis w.r.t. a sample $S$ is defined as ^$\hat{\mathcal{L}}(h,S)\triangleq\frac{1}{m}\sum_{j=1}^{m}\ell(h,z_{j})$.
In the following Section, and in Section 4, we first consider only two tasks from distributions $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$, referring to source and target. Within the classic PAC-Bayes setting (McAllester, 1999; Alquier, 2024), let $Q_{s}$ be a distribution over the set of hypotheses learned by some algorithm $J_{s}$ over $S_{s}\sim\mathcal{D}_{s}$ and a data-free prior hypothesis distribution $P$, such that $Q_{s}=J_{s}(S_{s},P).$ We then proceed to utilize another algorithm $J_{t}$ that operates on $S_{t}\sim\mathcal{D}_{t}$, making use of $Q_{s}$, such that $Q_{s:t}=J_{t}(S_{t},Q_{s}).$ For now, we make no assumptions on $J_{s}$ and $S_{t}$ other than their inputs. Of particular note is that $J_{t}$ has access to information on the previous task only via the prior distribution $Q_{s}$.
Definition 2.
The backwards transfer loss of $Q_{s:t}$ on task $\mathcal{D}_{s}$ is defined as
$\mathrm{BWT}(Q_{s:t},\mathcal{D}_{s})\triangleq\mathbb{E}_{h\sim Q_{s:t}}\left[\mathcal{L}(h,\mathcal{D}_{s})\right]=\mathcal{L}(Q_{s:t},\mathcal{D}_{s}).$
The forgetting of $Q_{s:t}$ on task $\mathcal{D}_{s}$ is defined as
$F(Q_{s:t},\mathcal{D}_{s})$ $\triangleq\mathrm{BWT}(Q_{s:t},\mathcal{D}_{s})-\mathbb{E}_{h\sim Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})\right]$
$=\mathcal{L}(Q_{s:t},\mathcal{D}_{s})-\mathcal{L}(Q_{s},\mathcal{D}_{s}).$
Intuitively, the backward transfer $\mathcal{L}(Q_{s:t},\mathcal{D}_{s})$ measures the performance of the updated model on the previous task $\mathcal{D}_{s}$, and the forgetting $F(Q_{s:t},\mathcal{D}_{s})$ measures how much worse this performance is compared to the loss immediately after learning $S_{s}\sim\mathcal{D}_{s}$. Since the continual learning setting assumes that tasks are given in order, when we are given task $\mathcal{D}_{t}$ we can no longer optimize $\mathcal{L}(Q_{s},\mathcal{D}_{s})$. As such, from an optimization perspective, the best we can do at any given point is to minimize the backwards transfer loss $\mathcal{L}(Q_{s:t},\mathcal{D}_{s})$. Note that our definition of forgetting is identical to that of Lin et al. (2023). Our definition of backward transfer could be equated to either the positive part of the forgetting loss or to their definition of “generalization error”, taken only on previous tasks.
We note that this definition refers to the test forgetting, meaning the model’s ability to still generalize well on previously seen domains, rather than measuring the retention of training performance on previous tasks. Due to this definition, simple measures such as memorization of previous training tasks cannot effectively minimize forgetting.
Definition 3.
The transfer loss of $Q_{s:t}$ on task $\mathcal{D}_{t}$ is defined as $\mathcal{L}(Q_{s:t},\mathcal{D}_{t})$, often referred to as the generalization loss.
Compared to the backward transfer, forward transfer (i.e., generalization) is better explored in general, and in the continual learning setting in particular, and several bounds are available (Bennani and Sugiyama, 2020; Benavides-Prado and Riddle, 2022).
Extending these notions to the full $T$ task setting, where tasks $\{1,\ldots,T\}$ appear sequentially, we define
$BWT(Q_{1:T},[T])\triangleq\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}(Q_{1:T},\mathcal{D}_{t})$
$F(Q_{1:T},[T])\triangleq\frac{1}{T-1}\sum_{t=1}^{T-1}\left(\mathcal{L}(Q_{1:T},\mathcal{D}_{t})-\mathcal{L}(Q_{1:t},\mathcal{D}_{t})\right)$
where $Q_{1:t}$ is the learned posterior based on $t$ tasks.
3 Data-dependent bounds for backward transfer
As we mentioned previously, in the continual learning setting we have no access to future tasks, and thus minimizing test forgetting reduces to minimizing the overall backward transfer as a proxy objective. In order to arrive at an upper bound on the backward transfer, we make use of concentration inequalities, relying on the change-of-measure inequality of Donsker and Varadhan (1975).
Theorem 3.1.
$\mathrm{(Forgetting)}$ For any fixed $S_{s},S_{t},Q_{s},Q_{s:t}$, and $\lambda_{t}>0$,
$\begin{split}&\mathcal{L}(Q_{s:t},\mathcal{D}_{s})\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right].\end{split}$ (1)
The full proof of Theorem 3.1 is provided in Appendix A. The main idea is to use Lemma A.1 (the change-of-measure inequality) with $f(z)=\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))$. Unlike standard (forward) transfer results, the l.h.s. depends on the source task $s$ while the r.h.s. depends on $t$. This is necessary since we no longer have access to the training data of the source task $S_{s}$. By subtracting $\mathcal{L}(Q_{s},\mathcal{D}_{s})$ from both sides, we arrive at an upper bound on the forgetting.
Looking at the three terms in the r.h.s. of (LABEL:eq:forget-base2) we see a complex tradeoff: if ^$\hat{\mathcal{L}}(Q_{s},S_{t})$ is small, the first two terms will be small as choosing $Q_{s:t}=Q_{s}$ or a posterior with low KL-divergence will lead to low empirical loss, but the final term will be larger (we subtract a smaller number in the exponent). If, on the other hand, ^$\hat{\mathcal{L}}(Q_{s},S_{t})$ is high, we can expect the disagreement term
$\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
to be small, but the first two terms will increase as achieving low ^$\hat{\mathcal{L}}(Q_{s:t},S_{t})$ requires higher KL-divergence. In both cases, if the prior $Q_{s}$ transfers well to the new task $S_{t}$, as expected for similar tasks, we see a lower r.h.s.
We note that unlike the other terms in the r.h.s. of (LABEL:eq:forget-base2), the expected loss on the previous task $\mathcal{L}(h,\mathcal{D}_{s})$ is not data-dependent. However, since this term does not depend on the current posterior $Q_{s:t}$, it has no direct effect on any optimization of the posterior.
Throughout the rest of the paper, we focus on classification problems, though our bounds hold for any setting where the conditions hold. Similarly to many PAC-Bayes bounds, results can be extended to unbounded losses (e.g. regression) under moment conditions (e.g. sub-Gaussian losses (Alquier et al., 2016; Alquier, 2024)).
Assumption 1.
For any hypothesis and data, the loss is bounded, $\ell(h,z)\in[0,K].$
The empirical loss can be removed from the final term in (LABEL:eq:forget-base2) leading to the following result.
Corollary 3.2.
Let $\Delta\mathcal{L}(h,s,t)\triangleq\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t})$. For any fixed $S_{s},Q_{s},\lambda_{t}>0$, the following applies uniformly for all posteriors $Q_{s:t}$ with probability at least $1-\delta$ over the choice of $S_{t}$ ($m_{t}=|S_{t}|$),
$\begin{split}\mathcal{L}(Q_{s:t},\mathcal{D}_{s})&\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right]\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log(1/\delta).\end{split}$ (2)
Proof of Corollary 3.2 is provided in Appendix A. The term $\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t})}\right]$ measures domain disagreement over $Q_{s}$, and can be hard to quantify in general.
For the $T$-task setting, we can bound the individual losses for each task. By using the same change of measure inequality as in Theorem 3.1, we can derive the following general bound on forgetting.
Theorem 3.3.
$\mathrm{(Forgetting)}$ For any $\lambda_{T}>0$, for any $S_{T}\sim\mathcal{D}_{T}$ and $t\in[T-1]$,
$\mathcal{L}$ $(Q_{1:T},\mathcal{D}_{t})\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})$
$+\frac{1}{\lambda_{T}}\log\mathbb{E}_{h\sim Q_{1:T-1}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{T}))}\right]$ (3)
By averaging Theorem 3.3 on all $T-1$ previous tasks, we get
Corollary 3.4.
For any fixed $Q_{1:T-1},Q_{1:T},S_{1},\ldots,S_{T}$, and $\lambda_{T}>0$,
$\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}(Q_{1:T},\mathcal{D}_{t})\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})$
$+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})$ (4)
$+\frac{1}{(T-1)\lambda_{T}}\sum_{t=1}^{T-1}\log\mathbb{E}_{Q_{1:T-1}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{T}))}\right].$
Subtracting the loss for each task from both sides and applying the same steps as Corollary 3.2, we get
Corollary 3.5.
($\mathrm{Forgetting}$) For any fixed $\lambda_{T}>0$ and $Q_{1:T-1},S_{1},\ldots,S_{T-1}$, the following applies uniformly for all posteriors $Q_{1:T}$ with probability at least $1-\delta$ over the choice of $S_{T}$,
$F(Q_{1:T},[T])\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})-\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}(Q_{1:t},\mathcal{D}_{t})$
$+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})+\frac{\lambda_{T}K^{2}}{8m_{T}}+\frac{\log{1/\delta}}{\lambda_{T}}$
$+\frac{1}{(T-1)\lambda_{T}}\sum_{t=1}^{T-1}\log\mathbb{E}_{Q_{1:T-1}}\left[e^{\lambda_{T}\Delta\mathcal{L}(h,T,t)}\right].$ (5)
4 Oracle bounds for forward and backward transfer
First, recall the Gibbs posterior in the single task setting.
Definition 4.
The empirical Gibbs posterior with parameter $\lambda$, is defined as
$\hat{Q}^{\lambda}_{s}(h)=\frac{P(h)e^{-\lambda\hat{\mathcal{L}}(h,S_{s})}}{\mathbb{E}_{h\sim P}\left[e^{\lambda\hat{\mathcal{L}}(h,S_{s})}\right]}~{}.$ (6)
While the general upper bounds on backward transfer in Section 3 are useful for designing theoretically motivated algorithms (see Section 5), there is merit in trying to better understand the behavior of these bounds for specific posterior distributions. To that end, we consider bounds on performance relative to that of an oracle who knows the data-distribution, as opposed to the data-dependent bounds established in Section 3. Specifically, we consider the continual Gibbs learner $\hat{Q}^{\lambda_{t}}_{s:t}(h)=\frac{Q_{s}(h)e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}}{\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}\right]}~{}.$ The Gibbs learner is of particular interest in the context of analyzing bounds with KL-divergence as it provides an explicit expression for the divergence for any prior. Note that as opposed to the more general Gibbs posterior of (6), here the distribution takes into account $S_{t}$ in addition to $S_{s}$.
Theorem 4.1.
For any $Q_{s},S_{s},\lambda_{t}>0$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}&\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\\ &\inf_{Q_{s:t}}\left\{\mathcal{L}(Q_{s:t},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right].\end{split}$ (7)
Proof of Theorem 4.1 is provided in Appendix A. The main idea of the proof is to choose the Gibbs learner as the posterior in (LABEL:eq:forget-base2). Equation (7) contains a domain disagreement term, $\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\triangleq\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right].$
Assuming tasks are sufficiently similar, we arrive at the following Corollary (proof in Appendix A).
Corollary 4.2.
For any given choice of $S_{s}$ and for any $\lambda_{t}>0$, if $\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\epsilon_{s,t}$, then $\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}F(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+2\epsilon_{s,t}.$
4.1 Oracle bounds with discrepancy terms
As we can see from Theorem 4.1, the behavior of $Q_{s}$ with regard to both tasks can affect the bound significantly. While Theorem 4.1 demonstrates the importance of $Q_{s}$, its exact effect is unclear. Next we discuss several corollaries that provide a clearer role for $Q_{s}$ that defines its desired behavior. The basic oracle inequality theorem we start from is the following (proof in appendix A).
Corollary 4.3.
For any given $S_{s}\sim\mathcal{D}_{s},Q_{s},\lambda_{t}>0$,
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}$ $\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}$
$+\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\mathcal{L}(h,\mathcal{D}_{s})\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right]},$ (8)
A more explicit bound can be derived from (4.3), see Corollary A.7 in the Appendix.
So far we have focused on oracle bounds for the Gibbs posterior for any prior $Q_{s}$. A further simplification can be obtained by considering the Gibbs prior (6), $Q_{s}=\hat{Q}^{\lambda_{t}}_{s}(h).$
From our definition of the domain disagreement term,
$\begin{split}\mathbb{E}_{S\sim\mathcal{D}_{s}}&\mathrm{Dis}(\hat{Q}^{\lambda_{t}}_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}-\mathcal{L}(P,\mathcal{D}_{t})\\ &-\frac{1}{\lambda_{t}}\mathbb{E}_{S\sim\mathcal{D}_{s}}\log\mathbb{E}_{h\sim P}\left[e^{-\lambda_{t}\hat{\mathcal{L}}(h,S)}\right].\end{split}$
Plugging this into (7) and using Jensen’s inequality, we obtain the following theorem.
Theorem 4.4.
For any $\lambda_{t}>0$, if $Q_{s}$ obeys (6),
$\begin{split}&\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s},\mathcal{D}_{t})\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\lambda_{t}K^{2}}{8m_{s}}+\mathcal{L}(P,\mathcal{D}_{s})-\mathcal{L}(P,\mathcal{D}_{t}).\end{split}$
This provides us with a more practical prior $\hat{Q}^{\lambda_{t}}_{s}$ at the cost of additional approximation error $\lambda_{t}K^{2}/8m_{s}$. If $m_{s},m_{t}\rightarrow\infty$ [^2] In this case, $\hat{Q}^{\lambda}_{s}=Q^{\lambda}_{s},\hat{Q}^{\lambda_{t}}_{s:t}=Q^{\lambda_{t}}_{s:t}$, we get an interesting bound involving generalization and forgetting,
$\begin{split}\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})&-\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s},\mathcal{D}_{t})\leq\\ &\mathcal{L}(P,\mathcal{D}_{s})-\mathcal{L}(P,\mathcal{D}_{t}).\end{split}$
We note that the right-hand-side is data-free and describes the general loss landscapes of both tasks with respect to the prior $P$. If both tasks come from the same distribution, for example, this implies that forgetting is upper bounded by the forward transfer (for Gibbs measures). This suggests that any problem with high forgetting for the Gibbs measure would also have poor forward transfer and a problem with good forward transfer will also have low forgetting.
4.2 Gibbs posteriors for correlated tasks
So far, we have derived oracle bounds that offer useful insights on the backward transfer for the Gibbs posterior. We would also like to examine whether stronger assumptions can offer bounds with improved or vanishing forgetting.
Theorem 4.5.
For any $\lambda_{t}>0$, if (6) holds, with $\mathcal{L}(P,\mathcal{D}_{s},\mathcal{D}_{t})\triangleq\mathcal{L}(P,\mathcal{D}_{s})+\mathcal{L}(P,\mathcal{D}_{t})$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\lambda_{t}K^{2}}{8m_{s}}$ (9)
$+\mathcal{L}(P,\mathcal{D}_{s},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim P}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{t})}\right],$
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\lambda_{t}K^{2}}{8m_{s}}$
$+\mathcal{L}(P,\mathcal{D}_{s},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim P}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{s})}\right].$
We can see that in both cases, the loss is bounded by functions of sample size originating from the deviations of sample mean from its true mean (e.g., Hoeffding’s inequality), and the loss of the data-free prior, which serves as a limit if no additional information on the tasks is available.
One such source of additional information is the loss covariance, measuring task similarity,
$\mathrm{cov}_{\lambda_{t}}(P,s,t)\triangleq\mathrm{cov}_{h\sim P}\left(e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{s})},e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}\right).$
We note that this covariance term is bounded in $[-1,1]$.
Corollary 4.6.
For any $\lambda>0$, if (6) holds, and $\mathrm{cov}_{\lambda_{t}}(P,s,t)\geq 0$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}+\mathcal{L}(P,\mathcal{D}_{s})$
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\mathcal{L}(P,\mathcal{D}_{t}).$
We can further improve upon this bound given a tighter bound on the covariance.
Corollary 4.7.
For any $\lambda>0$, if (6) holds, and $\mathrm{cov}_{\lambda_{t}}(P,s,t)\geq e^{-c}$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}+\frac{c}{\lambda_{t}}$
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{c}{\lambda_{t}}.$
Both Corollaries 4.6 and 4.7 are specific cases of more general Theorems that apply for any finite number of tasks. These bounds will be presented in the following subsection.
We observe that if the covariance is negative, Corollaries 4.6 and 4.7 do not hold, and we cannot improve on (9).This is unsurprising, as it means that the hypotheses that perform well on the source task perform poorly on the target and vice versa, which makes learning to generalize without forgetting impossible unless the initial prior is somehow already near-optimal. We note that the bound in Theorem 4.6 does not converge to zero as sample size goes to infinity, and the final term is in fact independent of the sample size. This is due to the fact that the covariance being non-negative is a necessary condition to guarantee that the backward transfer does not increase. As we can see in Corollary 4.7, a higher covariance leads to bounds that converge to zero upon appropriately choosing $\lambda_{t}$, as sample size increases.
Note that the phenomenon of self-forgetting seen in the NTK regime in (Karakida and Akaho, 2021) can also be seen here: if $m_{s}$ is small we can get forgetting even if the target task is identical (high covariance), and learning the same task again with new data may be worse than using all of the data as a single training set.
4.3 Extension to $T$ tasks
Suppose we are given a set of tasks $\{1,2,\ldots,T\}=[T]$, appearing sequentially. We would like to minimize the average backward transfer after all tasks, $BWT(Q_{1:T},[T])$. In general, this definition does not assume anything about the construction of the posterior, other than the fact that we have no access to samples from future (unseen) tasks while learning each specific posterior $Q_{1:t}$.
Starting from (3.3), we show in appendix A that if all of the priors are empirical Gibbs distributions, namely $\forall i\in\{2,\ldots,T\},~{}~{}\hat{Q}^{\lambda_{i}}_{1:i}(h)\propto\hat{Q}^{\lambda_{i-1}}_{1:i-1}(h)e^{-\lambda_{i}\hat{\mathcal{L}}(h,S_{i})}$, where $\hat{Q}^{\lambda_{1}}_{1:1}(h)\propto P(h)e^{-\lambda_{1}\hat{\mathcal{L}}(h,S_{1})}$, we have the following oracle bound.
Theorem 4.8.
$\mathrm{(Forgetting)}$ For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, and that $\mathrm{cov}_{P}(i,[T])\geq 0$, for any sample of training sets $S_{j}\sim\mathcal{D}_{j}$, $\forall i\in[T-1]$
$\begin{split}\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\mathcal{L}(P,\mathcal{D}_{i}).\end{split}$ (10)
The main ideas of the proof are to make use of the structure of Gibbs distributions to arrive at an explicit term for the KL-divergence, then make use of the assumption on the covariance to decompose said term into the loss on task $i$ and all other losses. As the extension to the covariance for two tasks, we have a similar notion of covariance
$\mathrm{cov}_{P}(i,[T])\triangleq\mathrm{cov}_{P}(e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})},e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}).$
As far as forward transfer, we can apply a similar analysis (for standard change-of-measure).
Theorem 4.9.
$\mathrm{(Transfer)}$ For any $\lambda_{T}>0$, if all $Q_{1:j}$ are empirical Gibbs posteriors, and $\mathrm{cov}_{P}(T,[T])\geq 0$, for any sample of training sets $S_{j<T}\sim\mathcal{D}_{j}$,
$\begin{split}\mathbb{E}_{S_{T}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{T})&\leq\frac{\lambda_{T}K^{2}}{8m_{T}}+\mathcal{L}(P,\mathcal{D}_{T}).\end{split}$ (11)
Theorem 4.8 is somewhat surprising, as it implies a sufficient condition for learning without forgetting that, for each individual task, does not become worse with the number of tasks and has no direct dependence on task order or on the length of time a task has not been seen. While the r.h.s. in (10) contains a constant term $\mathcal{L}(P,\mathcal{D}_{i})$, this term does not depend on the number of total tasks $T$ or on $i$. We also note that the forgetting $F(\hat{Q}^{\lambda_{T}}_{1:T},[T])$ may be negative. We show in Corollary A.9 that a stronger covariance condition even leads to vanishing forgetting, similarly to Corollary 4.7.
This lack of dependence on task order can be attributed to the nature of the empirical Gibbs distribution that applies a combination of exponential weights on the initial distribution $P$. The final weight of any given hypothesis $h$ therefore does not depend on the order of tasks but rather only on $P(h)$ and its empirical performance on each task.
The fact that this bound does not degrade with the number of tasks is a result of the assumption on the non-negative covariance: we assume that hypotheses that perform well on task $i$ tend to perform well on all other tasks, and thus the exponential weighting scheme does not significantly reduce their probability in the final distribution. While this is a somewhat strong assumption, it is weaker than assuming that a single optimal solution to all tasks exists.
5 Empirical validation
In this section we demonstrate the utility of our bounds on both synthetic and real-world data sets. We study three classes of task environments, with very different types of shifts between tasks, and show, that for sufficiently many tasks, the bounds are tight.
Since Corollary 3.4 applies uniformly over all posteriors, it can be optimized at each time step w.r.t. the current posterior. Specifically, we select
$\hat{Q}_{1:i}=\mathop{\mathrm{argmin}}_{Q_{1:i}}\left\{\hat{\mathcal{L}}(Q_{1:i},S_{i})+\frac{1}{\lambda_{i}}D_{\mathrm{KL}}(Q_{1:i}||Q_{1:i-1})\right\}$ (12)
This optimization process yields a practical continual learning algorithm that we can easily apply our upper bound on, detailed in Algorithm 1.
Algorithm 1 PAC-Bayes Continual Learning
1:  Continual-learn ($S_{1},\ldots,S_{T}$, $P$)
2:  Choose algorithmic parameters $\lambda_{1},\ldots,\lambda_{T}$
3:  Let $\hat{Q}_{1:0}(h)\triangleq P(h)$
4:  for each task $t$ from $1$ to $T$ do
5:     Update $\hat{Q}_{1:t}$ via (12), with $Q_{1:i-1}=\hat{Q}_{1:t-1}$
6:  end for
7:  Return $\hat{Q}_{1:T}$
This algorithm can be implemented in practice via variational inference (VI) using multivariate Gaussian prior and posterior distributions on model parameters. The posterior can be used to estimate the forgetting and its upper bound by using a pre-defined test set as a proxy for the data distribution $\mathcal{D}_{i}$. Calculating the upper bound (3.4) would add an additional calculation of $O(t\cdot(m+d))$ to each task, where $m=|S_{t}|$ and $d$ is the number of model parameters. In total, calculating the bound at each task will add $O(T^{2}\cdot(m+d))$. Compared to standard VI methods, the bound does not require additional memory or samples (it does require sampling from previous test tasks for calculating the bound, but this is also required for calculating forgetting in general).
Experimental settings
In order to measure the bound tightness of Algorithm 1, we must define the relevant metrics we wish to measure. These are the backwards transfer $BWT(\hat{Q}_{1:T},[T])$ (measured via a separate test set per task) and the forgetting $F(\hat{Q}_{1:T},[T])$, as well as the average forward transfer $\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}(\hat{Q}_{1:t},\mathcal{D}_{t})$. We compare Algorithm 1 to the expected maximum likelihood estimator (MLE) method, a baseline that only minimizes ^$\hat{\mathcal{L}}(Q_{1:t},S_{t})$ at each task $t$. Naturally, we would also like to consider the upper bound on both the forgetting and backward transfer.
In the following experiments, we use fully connected neural networks with a separate linear head per task to facilitate measuring the forgetting without re-training linear heads. We use Adam (Kingma and Ba, 2015) for optimization. The full list of hyper-parameters is listed in B.
10d Gaussian data
We begin by examining several setups of binary classification tasks in $\mathbb{R}^{10}$. All tasks draw samples from a $10$-dimensional Gaussian distribution $p(x)=\mathcal{N}(x;0,I_{10})$, and $y=\mathrm{sgn}(a^{\top}x)$. To simplify, only the first two features were used to determine $y$, meaning that there is a $2d$ linear separator embedded in $\mathbb{R}^{10}$. We consider the following settings: (1) Similar tasks: linear separators for all tasks are within $10^{\circ}$ of some reference angle. (2) Gradual shift: tasks change angle gradually in a set direction, with each consecutive task being within $10^{\circ}$ of the last one. (3) Orthogonal shift: Tasks come from two distinct, orthogonal angles, with the first half of tasks being from the first angle and the latter half corresponding to the second angle. Two additional settings are considered in Appendix B.
We used a total of $T=100$ tasks in this domain in order to observe general trends. We note that these problems differ significantly in task order and overall behavior, and are aimed to provide a diverse set of challenges. In addition, we should not expect a single algorithm to perform optimally across all settings, as known empirical methods based on regularization would likely perform poorly on orthogonal tasks whereas replay-based and architecture-based methods would have issues with gradual shift. We expect that in real-world settings, prior knowledge can be used to suggest specific algorithms. In particular, the settings of Gradual shift and Orthogonal shifts are constructed such that forgetting may be desirable. A recent paper by Kumar et al. (2023) discusses continual learning as a computationally constrained optimization problem and argues that forgetting non-recurring information is not “catastrophic”, especially given changing environments. Of our constructed settings, in both the gradual and orthogonal shift settings, not all previous knowledge is beneficial.
Table 1: Metrics For $10d$ Continual Tasks At $T=100$.
Method BWT BWT bound (3.4) Forgetting Forgetting bound (3.5) Test loss
Similar tasks
MLE $5.4\pm 0.5$ $7.2\pm 0.4$ $-10.9\pm 0.5$ $-9.1\pm 0.8$ $3.5\pm 0.3$
PAC-Bayes $\textbf{3.2}\pm 0.2$ $\textbf{5.3}\pm 0.3$ $-11.5\pm 1.0$ $-9.4\pm 1.1$ $\textbf{2.0}\pm 0.2$
Gradual shift
MLE $4.3\pm 0.3$ $6.4\pm 0.4$ $-10.1\pm 0.5$ $-8.0\pm 0.6$ $3.4\pm 0.2$
PAC-Bayes $\textbf{2.8}\pm 0.3$ $\textbf{4.8}\pm 0.3$ $-10.6\pm 0.8$ $-8.6\pm 0.8$ $\textbf{1.9}\pm 0.4$
Orthogonal shift
MLE $11.7\pm 2.7$ $13.9\pm 3.0$ $-6.8\pm 0.8$ $-4.6\pm 1.2$ $7.4\pm 2.5$
PAC-Bayes $\textbf{10.2}\pm 2.4$ $\textbf{12.0}\pm 2.3$ $-7.4\pm 1.3$ $-5.6\pm 1.8$ $\textbf{5.8}\pm 2.7$
(a) Gradual shift
(b) Orthogonal shift
Figure 1: Forgetting and upper bound (3.5) over time for $10d$ data. (a) Gradual shift (b) Task shift at $T=50$.
Table 1 describes the forward and backward transfer as well as forgetting for the $10d$ tasks. All test metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Backward transfer is the final loss percentage over all test sets. Standard error is reported after the $\pm$ sign. Best result for each setting is in bold. We can see that for most settings, Algorithm 1 tends to provide good forward transfer, as well as backward transfer and forgetting. For the orthogonal shift setting, we can see that adding a prior for regularization leads to looser upper bounds on both backward transfer and forgetting. Figure 1 compares the forgetting for the gradual shift setting compared to the more abrupt orthogonal shift at $T=50$. We can clearly see that the forgetting upper bound for the PAC-Bayesian method increases abruptly at the point of task shift, as both the KL-divergence and the discrepancy term increase. Due to the sudden shift, the model from the previous task serves as a poor prior and the model is forced to change significantly. We also note that across all settings, the PAC-Bayesian algorithm has lower standard-error and is more stable across different random seeds.
Notably, the final forgetting at $T=100$ as well as its upper bound are negative for all settings. The forgetting itself is negative and seems to follow a trend of increasing at the start, decreasing as general model performance improves and finally increasing towards zero forgetting. This trend is more pronounced in the orthogonal shift setting but occurs to a lesser extent in the other settings. The upper bound (3.4) decreases gradually starting from a positive value to a negative one. We observe that as the shared model improves, forgetting decreases, likely due to the learned shared representation improving overall across several tasks, and we see the forgetting trend towards zero.
An interesting observation is that the backward transfer (and forgetting) for the orthogonal shift setting is much higher. In this setting, forgetting may be a desired phenomenon, as the data distribution shifts in the middle of the learning process. We see that indeed this shift leads to higher forgetting. Since the MLE baseline has no regularization term, the change is significantly more abrupt.
In appendix B we also consider the discounted backward transfer and forgetting, using the formulae
$BWT^{\gamma}(Q_{1:T},[T])\triangleq\frac{1}{T-1}\sum_{t=1}^{T-1}\gamma^{T-1-t}\mathcal{L}(Q_{1:T},\mathcal{D}_{t}),$
$\begin{split}&F^{\gamma}(Q_{1:T},[T])\triangleq\\ &\frac{1}{T-1}\sum_{t=1}^{T-1}\gamma^{T-1-t}(\mathcal{L}(Q_{1:T},\mathcal{D}_{t})-\mathcal{L}(Q_{1:t}),\mathcal{D}_{t}).\end{split}$
These discounted forms allow us to re-weigh the forgetting based on recency, as achieving low forgetting on recent tasks may be more important in some settings than low overall forgetting. In general (for $\gamma=0.95$), the discounted backward transfer is lower, and the forgetting is higher. This is likely due to the shared model improving over time.
() BWT
() Forgetting
() Test loss
Figure 2: Metrics over time for the permuted-MNIST (top row) and split-MNIST (bottom row) tasksets. (a) backward transfer and its upper bound (3.4) (b) forgetting and its upper bound (3.5) (c) average test error over time.
Vision tasks
In this section, we examine Algorithm 1 on more realistic problem domains, namely permuted-MNIST (Goodfellow et al., 2015), a domain-incremental continual learning problem (De Lange et al., 2021) where a random permutation of the pixels is applied on each image for each task, and split-MNIST (Zenke et al., 2017), a sequential set of binary classification tasks constructed from the MNIST dataset. We used convolutional neural networks for these visual domains, and $T=80$ tasks in total for both.
Figure 2() shows the gradual change in backward transfer. We can see that after an initial warm-up period, both settings have gradually improving backward transfer as well as tighter upper bounds (3.3). Figure 2() shows that for both settings, the backward transfer loss behaves similarly to the test error, also decreasing with additional tasks. This suggests that the shared representation has sufficient network capacity to learn a useful representation for all MNIST vision tasks.
Figure 2() shows the forgetting and its upper bound (3.4) over time. For the split-MNIST problem, we see similar behavior to that of the $10d$ Gaussian data, where the forgetting increases, decreases and then begins to trend back upwards toward zero. For the permuted-MNIST problem we only see a general upward trend. The difference coincides well with the change in test loss (forward transfer) over time: in the permuted-MNIST setting we see constant improvement whereas whereas for the split-MNIST setting we see a plateau in forward transfer until roughly $T=20$ that corresponds to an increase in forgetting, followed by a decrease in both forgetting and forward transfer as the shared representation improves.
The full tables of the results appear in Appendix B
6 Conclusions
In this work, we derived several upper bounds on the test forgetting (via backward transfer) for both general model classes and for the Gibbs posterior, based on the change of measure approach. These upper bounds are data-dependent, potentially offering tighter bounds if improved prior models for the task are available, or if tasks are structured such that their loss landscapes are similar. In particular, we focused on oracle bounds for Gibbs posteriors that offered tight bounds on backward transfer if task losses are highly positively correlated, thus making the knowledge accumulation process highly effective for all tasks.
Based on our theoretical bounds, we constructed an algorithm for continual learning with empirically low forgetting and tight upper bounds. We examined this approach on several simple task constructions as well as slightly more complex vision tasks. In our experiments we noted a relation between forward and backward transfer, especially for mostly static settings such as the domain-incremental continual learning problem. As noted by several previous theoretical works and practical examinations (see Introduction), task order and similarity can greatly influence both forgetting and generalization. While this empirical demonstration is not the main focus of our paper, it suggests that weighting schemes based on notions of loss agreement merit further exploration for the domain incremental setting.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] General assumptions are clearly stated, specific assumptions are written for each theorem and corollary, and section $2$ details the overall setting.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] Algorithm 1 has a short analysis comparing its time and sample complexity properties compared to standard variational inference without the novel upper bound.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] Attached as a zip file as supplementary material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] General assumptions are clearly stated, specific assumptions are written for each theorem and corollary, and section $2$ details the overall setting.
    2. (b)
      Complete proofs of all theoretical results. [Yes] Appendix A contains full proofs of all theoretical results.
    3. (c)
      Clear explanations of any assumptions. [Yes] General assumptions are clearly stated, specific assumptions are written for each theorem and corollary.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] Code (with data download) is attached as a zip file, readme file contains the flags used for reproduction.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] detailed in Appendix B.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] As listed, standard error over $5$ random seeds is used for all measures.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] Detailed in Appendix B.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] Datasets are cited.
    2. (b)
      The license information of the assets, if applicable. [Yes] Dataset licence (GNU GPLv3) is included.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes] Code is included in the supplemental material.
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Appendix A Appendix - proofs
Lemma A.1.
Shui et al. (2020) Let $\pi$ and $\rho$ be two distributions on a common space $\mathcal{Z}$ such that $\rho$ is absolutely continuous w.r.t. $\pi$. For any $\lambda_{t}\in\mathbb{R}$ and any measurable function $f:\mathcal{Z}\rightarrow\mathbb{R}$ such that $\mathbb{E}_{z\sim\pi}\left[e^{\lambda_{t}(f(z)-\mathbb{E}_{\pi}f(z))}\right]<\infty$, we have
$\begin{split}\lambda_{t}&\left(\mathbb{E}_{z\sim\rho}\left[f(z)\right]-\mathbb{E}_{z\sim\pi}\left[f(z)\right]\right)\leq D_{\mathrm{KL}}(\rho||\pi)+\log\mathbb{E}_{z\sim\pi}\left[e^{\lambda_{t}(f(z)-\mathbb{E}_{\pi}f(z))}\right],\end{split}$ (13)
where $D_{\mathrm{KL}}$ is the KL-divergence and equality is achieved for $f(z)=\mathbb{E}_{z\sim\pi}f(z)+\frac{1}{\lambda_{t}}\log(\frac{d\rho}{d\pi})$.
Theorem A.2.
Restatement of Theorem 3.1: For any fixed $S_{s},S_{t},Q_{s},Q_{s:t}$, for any $\lambda_{t}>0$,
$\begin{split}\mathcal{L}(Q_{s:t},\mathcal{D}_{s})&\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]\end{split}$ (14)
Proof.
Starting from Lemma A.1, we can choose $f(z)=\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))$, giving us
$\lambda_{t}\mathbb{E}_{h\sim Q_{s:t}}\left[\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t})\right]-\lambda_{t}\mathbb{E}_{h\sim Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t})\right]$
$~{}~{}\leq D_{\mathrm{KL}}(Q_{s:t}||Q_{s})+\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}e^{-\lambda_{t}(\mathcal{L}(Q_{s},\mathcal{D}_{s})-\hat{\mathcal{L}}(Q_{s},S_{t}))}\right]$
Extracting terms that do not depend on $h$ from the expectation, we get
$\begin{split}F(Q_{s:t},\mathcal{D}_{s})&\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})-\mathcal{L}(Q_{s},D_{s})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]\end{split}$ (15)
Lemma A.3.
(See also Lemma $1.1$ in Alquier (2024)) Let $l:Z\times H\rightarrow[0,K]$ be a measurable function. Let $\pi\in\mathcal{M}(H)$ be a distribution over $H$ that is independent w.r.t. $Z$. Let $S\in Z^{m}$ be an i.i.d. sample. With probability at least $1-\delta$ over the choice of $S$,
$\log\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]\leq\frac{t^{2}K^{2}}{8m}+\log{1/\delta}$
Proof.
Using Markov’s inequality, we know that
$\textrm{Pr}\left(\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]<\frac{1}{\delta}\mathbb{E}_{S\sim Z^{m}}\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]\right)\geq 1-\delta$
Applying Fubini’s theorem (both distributions are independent), we can re-order the expectations
$\textrm{Pr}\left(\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]<\frac{1}{\delta}\mathbb{E}_{h\sim\pi}\mathbb{E}_{S\sim Z^{m}}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]\right)\geq 1-\delta$
Since $S$ is drawn i.i.d. and $l$ is bounded, we can apply Hoeffding’s lemma to each example, giving us
$\textrm{Pr}\left(\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]<\frac{1}{\delta}\mathbb{E}_{h\sim\pi}\left[e^{\frac{t^{2}K^{2}}{8m}}\right]\right)\geq 1-\delta$
$\textrm{Pr}\left(\log\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]<\log\frac{1}{\delta}e^{\frac{t^{2}K^{2}}{8m}}\right)\geq 1-\delta$
and so we have
$\textrm{Pr}\left(\log\mathbb{E}_{h\sim\pi}\left[e^{t(\frac{1}{m}\sum_{i}l(z_{i},h)-\mathbb{E}_{z}l(z,h))}\right]<\log\frac{1}{\delta}+\frac{t^{2}K^{2}}{8m}\right)\geq 1-\delta$
Corollary A.4.
Restatement of Corollary 3.2: If $\ell\in[0,K]$, for any fixed $S_{s},Q_{s},\lambda_{t}>0$, the following applies uniformly for all posteriors $Q_{s:t}$ with probability at least $1-\delta$ over the choice of $S_{t}$,
$\begin{split}\mathcal{L}(Q_{s:t},\mathcal{D}_{s})&\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\\ &+\frac{1}{2\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]+\frac{\lambda_{t}K^{2}}{4m_{t}}+\frac{1}{2\lambda_{t}}\log(1/\delta)\end{split}$ (16)
Proof.
Starting from (15), we note that
$\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]=\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t})+\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right].$
This gives us
$\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]$ $=\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
$\triangleq\log\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}e^{\lambda_{t}\Delta\hat{\mathcal{L}}(h,\mathcal{D}_{t},S_{t})}\right].$
Similarly to Lemma A.3 we can apply Fubini’s theorem, and since $\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})$ is independent of $S_{t}$ we can re-order the expectations and use Hoeffding’s lemma on $e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{t},S_{t})}$ if $\ell\in[0,K]$. More formally,
$\mathbb{E}_{S_{t}}\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}e^{\lambda_{t}\Delta\hat{\mathcal{L}}(h,\mathcal{D}_{t},S_{t})}\right]=\mathbb{E}_{Q_{s}}\mathbb{E}_{S_{t}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}e^{\lambda_{t}\Delta\hat{\mathcal{L}}(h,\mathcal{D}_{t},S_{t})}\right]$
$=\mathbb{E}_{Q_{s}}\left[\mathbb{E}_{S_{t}}\left[e^{\lambda_{t}\Delta\hat{\mathcal{L}}(h,\mathcal{D}_{t},S_{t})}\right]e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]\leq\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}^{2}K^{2}/8m_{t}}e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]$
Finally, we get with probability at least $1-\delta$ over the choice of $S_{t}$
$\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]\leq\log\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]+\frac{\lambda_{t}^{2}K^{2}}{8m_{t}}+\log(1/\delta)$
Theorem A.5.
Restatement of Theorem 4.1: For any $Q_{s},S_{s},\lambda_{t}>0$, if $\ell\in[0,K]$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}&\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\inf_{Q_{s:t}}\left\{\mathcal{L}(Q_{s:t},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]\end{split}$ (17)
Proof.
Starting from Lemma A.1, we know that
$\mathbb{E}_{z\sim\rho}\left[f(z)\right]\leq\mathbb{E}_{z\sim\pi}\left[f(z)\right]+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(\rho||\pi)+\frac{1}{\lambda_{t}}\log\mathbb{E}_{z\sim\pi}\left[e^{\lambda_{t}(f(z)-\mathbb{E}_{\pi}f(z))}\right]$
In particular, for $\hat{\rho}_{\lambda}(z)\propto\pi(z)e^{-\lambda_{t}f(z)}$, this is an equality (from Donsker and Varadhan’s [1975] variational lemma). From this, we know that
$\mathbb{E}_{z\sim\hat{\rho}_{\lambda}}\left[f(z)\right]=\mathbb{E}_{z\sim\pi}\left[f(z)\right]+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(\hat{\rho}_{\lambda}||\pi)+\frac{1}{\lambda_{t}}\log\mathbb{E}_{z\sim\pi}\left[e^{\lambda_{t}(f(z)-\mathbb{E}_{\pi}f(z))}\right]$ (18)
If we pick $f(z)=\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t})$ as before, we get
$\begin{split}\mathbb{E}_{z\sim\hat{\rho}_{\lambda}}\left[\mathcal{L}(z,\mathcal{D}_{s})\right]&=\mathbb{E}_{z\sim\pi}\left[f(z)\right]+\mathbb{E}_{z\sim\hat{\rho}_{\lambda}}\left[\hat{\mathcal{L}}(z,S_{t})\right]+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(\hat{\rho}_{\lambda}||\pi)\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{z\sim\pi}\left[e^{\lambda_{t}(f(z)-\mathbb{E}_{\pi}f(z))}\right]\end{split}$
And as such,
$F(\hat{\rho}_{\lambda},\mathcal{D}_{s})\leq\inf_{\rho}\left\{\hat{\mathcal{L}}(\rho,S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(\rho||\pi)\right\}-\mathcal{L}(\pi,D_{s})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{z\sim\pi}\left[e^{\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))}\right],$
or using our previous terminology with $\hat{Q}^{\lambda_{t}}_{s:t}(h)\propto Q_{s}(h)e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}$,
$\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\inf_{Q_{s:t}}\left\{\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
If we take an expectation on $S_{t}$, we get
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})&\leq\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\inf_{Q_{s:t}}\left\{\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}\\ &+\frac{1}{\lambda_{t}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]\end{split}$
This gives us the following oracle inequality (in expectation):
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})&\leq\inf_{Q_{s:t}}\left\{\mathcal{L}(Q_{s:t},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]\end{split}$
We have
$\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
$=\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
$=\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
Using Hoeffding’s lemma, we get
$\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right]\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]$
Proof of accompanying Corollary:
Proof.
From our assumption $0\leq\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\epsilon_{s,t}$ we also have $0\leq|\mathcal{L}(Q_{s},\mathcal{D}_{s})-\mathcal{L}(Q_{s},\mathcal{D}_{t})|\leq\epsilon_{s,t}$ since
$\mathbb{E}_{h\sim Q_{s}}\left[\Delta\mathcal{L}(h,s,t)\right]\leq\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t}),\quad\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\max_{h}\left[\Delta\mathcal{L}(h,s,t)\right].$
Looking at the negative transfer, we see
$\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}F(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\left[\mathcal{L}(Q_{s},\mathcal{D}_{t})-\mathcal{L}(Q_{s},\mathcal{D}_{s})\right]+\frac{\lambda_{t}K^{2}}{8m_{t}}+\epsilon_{s,t}$
and similarly from our assumption, we have
$\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}F(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+2\epsilon_{s,t}$
Corollary A.6.
Restatement of Corollary 4.3: If $\ell\in[0,K]$, for any given $S_{s}\sim\mathcal{D}_{s},Q_{s},\lambda_{t}>0$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}&\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\mathcal{L}(h,\mathcal{D}_{s})\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}\end{split}$ (19)
Proof.
We can derive such a bound using (7) by setting the optimal posterior:
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})&\leq-\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{t})}\right]\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]\end{split}$
This is the same as writing
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{t})}\right]}$
Since $e^{k}\geq 0$ for all $k\in\mathbb{R}$, we can apply the log-sum inequality:
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\log\frac{e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}}{e^{\lambda_{t}(-\mathcal{L}(h,\mathcal{D}_{t}))}}\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}$
Corollary A.7.
Appendix-only Corollary: If $\ell\in[0,K]$, for any given $Q_{s},\lambda_{t}>0,S_{s}\sim\mathcal{D}_{s}$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},&\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}\\ &+\frac{\sqrt{\mathbb{E}_{Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}}{\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}\cdot\left(\sqrt{\mathrm{Var}_{Q_{s}}(\mathcal{L}(h,\mathcal{D}_{s}))}+\mathcal{L}(Q_{s},\mathcal{D}_{s})\right)\end{split}$ (20)
Proof.
Starting from (4.3), we can apply the Cauchy-Schwartz theorem on the expectation and get
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\sqrt{\mathbb{E}_{Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]\mathbb{E}_{Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})^{2}\right]}}{\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]};$
$=\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\sqrt{\mathbb{E}_{Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}}{\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}\sqrt{\mathbb{E}_{Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})^{2}\right]}$ (21)
And using the definition of variance and the triangle inequality (since both terms are non-negative), we get (20).
Theorem A.8.
Restatement of Theorem 4.8: For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, $\ell\in[0,K]$, and that $\mathrm{cov}_{P}(e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})},e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})})\geq 0$, for any sample of training sets $S_{j}\sim\mathcal{D}_{j}$,
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\mathcal{L}(P,\mathcal{D}_{i})\end{split}$ (22)
Proof.
Starting from (3.3), suppose we assume that for each task we apply an empirical Gibbs learner, meaning
$\forall i\in[2,T],\hat{Q}^{\lambda_{i}}_{1:i}(h)\propto\hat{Q}^{\lambda_{i-1}}_{1:i-1}(h)e^{-\lambda_{i}\hat{\mathcal{L}}(h,S_{i})}$
where $\hat{Q}^{\lambda_{1}}_{1:1}(h)\propto P(h)e^{-\lambda_{1}\hat{\mathcal{L}}(h,S_{1})}$.
We can provide bounds on the forgetting of $\hat{Q}^{\lambda_{T}}_{1:T}$ by setting the posterior $Q_{1:T}$ and the prior $Q_{1:T-1}$ as Gibbs posteriors:
$\begin{split}\forall i\in[1,T-1]:\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{1}{\lambda_{T}}\log\frac{\mathbb{E}_{h\sim\hat{Q}^{\lambda_{T-1}}_{1:T-1}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{i})-\hat{\mathcal{L}}(h,S_{T}))}\right]}{\mathbb{E}_{h\sim\hat{Q}^{\lambda_{T-1}}_{1:T-1}}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{T})}\right]}\end{split}$
We can unravel the expectations and arrive at:
$\begin{split}\forall i\in[1,T-1]:\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{1}{\lambda_{T}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{\lambda_{T}\mathcal{L}(h,\mathcal{D}_{i})-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}\end{split}$
Taking an expectation over $S_{i}$,
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{1}{\lambda_{T}}\mathbb{E}_{S_{i}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{\lambda_{T}\mathcal{L}(h,\mathcal{D}_{i})-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}\end{split}$
Applying Jensen’s inequality:
$\begin{split}\forall i\in[1,T-1]:&\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})\\ &\leq\frac{1}{\lambda_{T}}\mathbb{E}_{S_{i}}\log\frac{\mathbb{E}_{h\sim P}\left[\mathbb{E}_{S_{i}}\left[e^{\lambda_{T}\mathcal{L}(h,\mathcal{D}_{i})-\lambda_{i}\hat{\mathcal{L}}(h,S_{i})}\right]e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}\end{split}$
If $\lambda_{i}=\lambda_{T}$, we can apply Hoeffding’s lemma:
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{i}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}\end{split}$
As in the paper, we mark
$\mathrm{cov}_{P}(i,[T])\triangleq\mathrm{cov}_{P}(e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})},e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}).$
From the definition of covariance, we can decompose the denominator:
$\begin{split}&\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})\leq\frac{\lambda_{T}K^{2}}{8m_{i}}\\ &+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{i}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right]+\mathrm{cov}_{P}(i,[T])}\end{split}$ (23)
And if this covariance is non-negative, we can set it at 00 and still retain a valid upper bound, giving us the forgetting bound via Jensen’s inequality.
Similarly for forward transfer, we get
$\begin{split}&\mathbb{E}_{S_{T}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{T})\leq\frac{\lambda_{T}K^{2}}{8m_{T}}\\ &+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{T}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T-1}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T-1}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{T})}\right]+\mathrm{cov}_{P}(T,[T])}\end{split}$ (24)
and apply $\mathrm{cov}_{P}(T,[T])\geq 0$ and Jensen’s inequality to get the bound for generalization.
Corollary A.9.
: Appendix-only Corollary: Under the same conditions as Theorem 4.8, if we additionally have
$\begin{split}\mathrm{cov}_{P}(e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})},e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})})\geq e^{-c}-\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right],\end{split}$
we have (for any sample of training sets $S_{j}\sim\mathcal{D}_{j}$),
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{c}{\lambda_{T}}.\end{split}$
Proof.
Considering our forgetting bound (23), we consider when
$\begin{split}\frac{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right]+\mathrm{cov}_{P}(i,[T])}\leq e^{c},\end{split}$
Moving terms around, this condition is satisfied if
$\begin{split}\frac{\mathrm{cov}_{P}(i,[T])}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}+\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right]\geq e^{-c},\end{split}$
and we can further simplify this by taking a slightly looser condition:
$\begin{split}\mathrm{cov}_{P}(i,[T])\geq e^{-c}-\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right]\end{split}$
and this is the condition we assumed. As such, we can replace the last term in (23) with
$\frac{1}{\lambda_{t}}\log e^{c},$
giving us a bound of the form:
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{c}{\lambda_{T}}.\end{split}$ (25)
Appendix B Appendix - configuration and hyper-parameters
For $10d$-Gaussian tasks, we use several setups of binary classification tasks in $\mathbb{R}^{10}$. All tasks draw samples from a $10$-dimensional Gaussian distribution $p(x)=\mathcal{N}(x;0,I_{10})$, and $y=\mathrm{sgn}(a^{\top}x)$. We consider several settings for $a$, where for all setting, all values of $a$ other than the first two are zero, meaning the last $8$ features do not impact the label. We consider the following settings:
  • Similar tasks: linear separators for all tasks are within $10^{\circ}$ of some reference angle.
  • Gradual shift: tasks change angle gradually in a set direction, with each consecutive task being within $10^{\circ}$ of the last one.
  • Orthogonal shift: two sets of similar tasks with a $90^{\circ}$ angle between the separators of both task sets. The first half of tasks are of the first type and the second half corresponds to the second type.
For $10d$-Gaussian tasks, we use $2000$ samples per task, and a total of $T=100$ tasks.
Table 2: Discounted ($\gamma=0.95$) backward transfer and forgetting for $10d$ continual tasks at $T=100$. Metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Standard error is reported after the $\pm$ sign.
Method Discount BWT Discount BWT bound Discount forget. Discount forget. bound
Similar tasks
MLE $0.7\pm 0.0$ $2.3\pm 0.1$ $-0.4\pm 0.0$ $1.2\pm 0.1$
PAC-Bayes $0.4\pm 0.0$ $2.1\pm 0.0$ $-0.4\pm 0.0$ $1.3\pm 0.1$
Gradual shift
MLE $0.6\pm 0.0$ $2.3\pm 0.1$ $-0.3\pm 0.0$ $1.3\pm 0.1$
PAC-Bayes $0.4\pm 0.0$ $2.0\pm 0.1$ $-0.3\pm 0.1$ $1.3\pm 0.1$
Orthogonal shift
MLE $1.6\pm 0.6$ $3.2\pm 0.6$ $-0.5\pm 0.1$ $1.1\pm 0.2$
PAC-Bayes $1.3\pm 0.6$ $2.9\pm 0.5$ $-0.6\pm 0.2$ $1.0\pm 0.3$
The model consists of a shared fully connected layer of $64$ neurons and a tanh activation, as well as a linear classification head for each task. Each task was trained for $1$ epochs with batch size $64$. The learning rate was static at $1e^{-3}$. The $\lambda$ parameter was set to $1e^{-2}$. All results were run for $5$ random seeds and averages and standard error were reported.
For permuted-MNIST, we used a different pixel permutation per task, and each task involved $10$-way classification. All $60000$ training samples were used for training with a batch size of $128$. The learning rate was static at $1e^{-3}$ and the $\lambda$ parameter was set to $1e^{-3}$. The model consists of a shared convolutional neural network consisting of 2 CNN blocks each comprised of $64$ $5\times 5$ $2d$ convolutions followed by max-pooling and tanh non-linearity, followed by a linear layer consisting of $400$ neurons and a tanh activation as well as a linear classification head for each task. All results were run for $5$ random seeds and averages and standard error were reported in Table 3.
Table 3: Metrics for permuted-MNIST tasks. All test metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Backward transfer is the final loss percentage over all test sets. Standard error is reported after the $\pm$ sign.
Method BWT BWT bound (3.4) Forgetting Forgetting bound (3.5) Test loss
MLE $21.1\pm 0.1$ $23.4\pm 0.4$ $-6.9\pm 0.2$ $-4.7\pm 0.4$ $12.4\pm 0.5$
PAC-Bayes $21.2\pm 0.2$ $22.8\pm 0.4$ $-6.9\pm 0.2$ $-5.2\pm 0.3$ $12.9\pm 0.3$
For split-MNIST, each task involved half of the labels (at random) chosen as positive and half as negative. This is a minor departure from the standard split-MNIST problem where $5$ different binary classification tasks are created and their loss is averaged, but the overall behavior is similar. All $60000$ training samples were used for training with a batch size of $256$. The learning rate was static at $1e^{-3}$ and the $\lambda$ parameter was set to $1e^{-4}$. The model used is identical to the CNN used for permuted-MNIST. All results were run for $5$ random seeds and averages and standard error were reported in Table 4.
Table 4: Metrics for split-MNIST tasks. All test metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Backward transfer is the final loss percentage over all test sets. Standard error is reported after the $\pm$ sign.
Method BWT BWT bound (3.4) Forgetting Forgetting bound (3.5) Test loss
MLE $19.0\pm 1.4$ $21.8\pm 1.2$ $-10.4\pm 0.4$ $-7.6\pm 0.6$ $12.8\pm 2.4$
PAC-Bayes $18.4\pm 1.1$ $21.9\pm 1.2$ $-10.2\pm 0.4$ $-6.8\pm 0.6$ $11.6\pm 1.9$
All experiments were run on local hardware with an NVIDIA GeForce 1080 GPU and a quad-core Intel i7 CPU.